OpenStudy (mrnebula):

1. Simplify by Dividing (8m^7) - (10m^7) / (2m^3)

OpenStudy (mrnebula):

@welshfella

OpenStudy (mrnebula):

This is what I got 8m^7 - 5m^4

OpenStudy (astrophysics):

\[\frac{ (8m^7)-(10m^7) }{ 2m^3 } \implies \frac{ 8m^7 }{ 2m^3 }-\frac{ 10m^7 }{ 2m^3 }\]

OpenStudy (astrophysics):

\[\frac{ x^m }{ x^n } = x^{m-n}\]

OpenStudy (mrnebula):

Ah ok, and thanks for the formula. :D